# Knowledge Distillation-Based Model Extraction Attack using Private Counterfactual Explanations

### v1

@@ Abstract
In recent years, there has been a notable increase in the deployment of machine learning (ML) models as services (MLaaS) across diverse production software applications. In parallel, explainable AI (XAI) continues to evolve, addressing the necessity for transparency and trustworthiness in ML models. XAI techniques aim to enhance the transparency of ML models by providing insights, in terms of model’s explanations, into their decision-making process. Simultaneously, some MLaaS platforms now offer explanations alongside the ML prediction outputs. This setup has elevated concerns regarding vulnerabilities in MLaaS, particularly in relation to privacy leakage attacks such as model extraction attacks (MEA). This is due to the fact that explanations can unveil insights about the inner workings of the model which could be exploited by malicious users. In this work, we focus on investigating how model explanations, particularly Generative adversarial networks (GANs)-based counterfactual explanations (CFs), can be exploited for performing MEA within the MLaaS platform. We also delve into assessing the effectiveness of incorporating differential privacy (DP) as a mitigation strategy. To this end, we first propose a novel MEA methodology based on Knowledge Distillation (KD) to enhance the efficiency of extracting a substitute model of a target model exploiting CFs. Then, we advise an approach for training CF generators incorporating DP to generate private CFs. We conduct thorough experimental evaluations on real-world datasets and demonstrate that our proposed KD-based MEA can yield a high-fidelity substitute model with reduced queries with respect to baseline approaches. Furthermore, our findings reveal that the inclusion of a privacy layer impacts the performance of the explainer, the quality of CFs, and results in a reduction in the MEA performance.
Keywords Explainable Artificial Intelligence  $\cdot$ Knowledge Distillation  $\cdot$ Privacy
@@ 1 Introduction
Numerous platforms currently offer MLaaS such as Google Cloud, Microsoft, IBM, and Amazon Web Services [9, 6, 10, 8, 7, 1]. The API’s input-output format for these services is publicly accessible, meaning that users have knowledge of the format of input data required by the service (more specifically, by the ML model), and can interpret its outputs. This opens the door for malicious users (or, more precisely, attackers) to attempt to perform privacy-breaching attacks such as membership inference attacks (MIA) [47], model inversion attacks (MINA) [51], and model extraction attacks (MEA) [49]. For instance, in MIA, an attacker aims to infer the presence of a specific individual instance in the private training set. In MINA, the attacker attempts to reconstruct the training set. In MEA, the attacker attempts to extract the model by training a substitute model using labeled data acquired through repetitive queries to the service provider’s confidential model. Several mitigation strategies for these security- and privacy-breaching attacks are already employed [17]. However, these strategies are faced with new challenges as emerging transparency requirements demand explanations for ML model decisions. In fact, MLaaS platforms are already shifting towards providing explanations alongside the predictions made by the deployed ML models [10, 7, 1, 8]. The need to provide users with statistics such as model explanations is linked to providing users with a transparent decision-making process of data-driven systems. The urge to enhance transparency through model explanations aligns, on one hand, with the aim of meeting recent regulatory (and compliance) demands, and on the other hand, with the aim of providing users with actionable insights. Such explanations are extracted using explainable AI (XAI) techniques, which aim to provide insight into how a model arrived at its decision [20, 43]. In this context, explanations may inadvertently provide attackers with insights to enhance their attacks, posing a fresh obstacle to existing security measures.
@@
In this work, we focus on investigating how example-based explanations, specifically counterfactual (CF) explanations, can be exploited to perform MEA within MLaaS scenarios. We precisely focus on CFs due to the unique insights they provide, as they reveal how to minimally alter original data instances to achieve a different model’s outcome, allowing users to interpret the dynamics underlying the model’s prediction shift from the decision boundary [15]. However, due to the proximity of CFs to the decision boundary and their representativeness of the training data, CFs possess a dual nature, serving not only as interpretative tools but also as potential assets for enhancing attacks as they provide additional insights and knowledge for potentially exploiting the vulnerabilities of the model under attack. Specifically, we propose a novel approach based on the concept of knowledge distillation (KD) to perform MEA with CFs as the representation of data. The unique capabilities of KD allow us to handle the problem not just as a standard classification task but as an estimation of the output probability distribution. Moreover, we propose a mitigation approach that employs differential privacy (DP) within the CF generator training pipeline, aiming to reduce the risk associated with providing CF explanations. Our analysis examines the attack performance of our proposed method in comparison to baselines. Additionally, we assess the impact of DP on the quality of GAN-based CFs and the performance of the MEA. The contribution of this paper is summarized as follows:
@@
  • KD-Based MEA with CFs: This research investigates potential vulnerabilities in MLaaS, with a specific focus on MEA facilitated by KD as extraction techniques and CFs as representatives of the training set. The research simulates an adversarial scenario wherein an attacker employs our proposed KD techniques to extract a substitute version of a target model.
@@ 2 Related Work
Several works have proposed strategies for performing MEA on classification and text generation tasks. Authors in [49] perform successful MEA on different ML models like decision trees, Support Vector Machines, and DNNs by using equation-solving and path-finding algorithms and learning theory. In [39] authors perform MEA on DNNs using active learning with unannotated public data. In [25], authors perform MEA in natural language processing on bert-based APIs, in which they explore transfer learning for MEA. Other works have focused on proposing querying strategies for MLaaS to effectively query the target model and extract accurate insights in addition to performing MEA [38, 23, 36]. In [38] authors train a knockoff network with queried image predictions and propose a reinforcement learning approach that additionally improves query sample efficiency in certain settings and provides performance gains. In [23] authors propose a new method that generates synthetic queries and optimizes training hyperparameters, and then propose a method to detect generic and effective detection of DNN MEA. In [36] authors investigate the type and amount of internal information about the black-box model that can be extracted from querying the deployed model, such as the architecture, optimization procedure, or training data. Authors in [21] improves the query efficiency of attacks by designing learning-based methods and focusing on the real accuracy of the extracted model.
@@
Similar to these works, our work aligns with exploring how CF explanations can be exploited for performing MEA. However, our contribution lies in modeling a novel approach to MEAs that exploits more effectively the insights that CFs carry with respect to other approaches that directly train a substitute model on extracted CFs. Moreover, in contrast to previous studies [12, 52], our work presumes that the attacker possesses no prior knowledge of the training set, leading them to query the model with zero knowledge.
@@
In terms of mitigation strategies, few works have focused on proposing methodologies to generate explanations while mitigating the revelation of sensitive insights about the decision boundary, the training set, or the model architectures. For instance, authors in [40] propose DP algorithms to construct feature-based model explanations, while authors in [54] propose to generate differentially private CFs via functional mechanisms. [41] proposes an approach for generating recourse paths leveraging differentially private clustering and demonstrates that constructing a graph on the cluster centers provides private recourse paths as CFs.
@@
Also, our work employs DP, however, by integrating it into the CF generation process, and eliminates the necessity for a distinct method to generate private CFs. We leverage the strength of generative adversarial networks (GANs)-based generators in producing high-quality, yet private, CFs. To the best of our knowledge, our work is the first to adapt DP within CF generative-based methods. We specifically focus on GANs-based CF generators due to their ability to generate high-quality CFs compared to other CF generation methods. We show the robustness of the DP-based proposed method against MEA, and we provide a detailed analysis that sheds light on the impact of CFs in exposing the distribution of the training set through well-defined metrics.
3 Preliminaries

### v2

# Knowledge Distillation-Based Model Extraction Attack using GAN-based Private Counterfactual Explanations

@@ Abstract
In recent years, there has been a notable increase in the deployment of machine learning (ML) models as services (MLaaS) across diverse production software applications. In parallel, explainable AI (XAI) continues to evolve, addressing the necessity for transparency and trustworthiness in ML models. XAI techniques aim to enhance the transparency of ML models by providing insights, in terms of model’s explanations, into their decision-making process. Simultaneously, some MLaaS platforms now offer explanations alongside the ML prediction outputs. This setup has elevated concerns regarding vulnerabilities in MLaaS, particularly in relation to privacy leakage attacks such as model extraction attacks (MEA). This is due to the fact that explanations can unveil insights about the inner workings of the model which could be exploited by malicious users. In this work, we focus on investigating how model explanations, particularly counterfactual explanations (CFs), can be exploited for performing MEA within the MLaaS platform. We also delve into assessing the effectiveness of incorporating differential privacy (DP) as a mitigation strategy. To this end, we first propose a novel approach for MEA based on Knowledge Distillation (KD) to enhance the efficiency of extracting a substitute model of a target model exploiting CFs, without any knowledge about the training data distribution by the attacker. Then, we advise an approach for training CF generators incorporating DP to generate private CFs. We conduct thorough experimental evaluations on real-world datasets and demonstrate that our proposed KD-based MEA can yield a high-fidelity substitute model with a reduced number of queries with respect to baseline approaches. Furthermore, our findings reveal that including a privacy layer can allow mitigating the MEA. However, on the account of the quality of CFs, impacts the performance of the explanations.
@@ 1 Introduction
Numerous platforms offer MLaaS such as Google Cloud, Microsoft, IBM, and Amazon Web Services [9, 6, 10, 8, 7, 1]. The API’s input-output format for these services is publicly accessible, meaning that users have knowledge of the format of input data required by the service (by the ML model), and can interpret its outputs. This opens the door for malicious users (or, more precisely, attackers) to attempt to perform privacy-breaching attacks such as membership inference attacks (MIA) [54], model inversion attacks (MINA) [59], and model extraction attacks (MEA) [57]. For instance, in MIA, an attacker aims to infer the presence of a specific individual instance in the private training set. In MINA, the attacker attempts to reconstruct the training set. In MEA, the attacker attempts to extract the model by training a substitute model using data acquired through repetitive queries to the service provider’s confidential model. Several mitigation strategies for these security- and privacy-breaching attacks are already employed [17]. However, these strategies are faced with new challenges as emerging transparency requirements demand explanations for ML model decisions. In fact, MLaaS platforms are shifting towards providing explanations alongside the predictions made by the deployed ML models [10, 7, 1, 8]. The need to provide users with such model explanations is linked to providing users with a transparent decision-making process of data-driven systems [40]. The urge to enhance transparency through model explanations aligns, on one hand, with the aim of meeting recent regulatory (and compliance) demands, and on the other hand, with the aim of providing users with actionable insights [38, 56, 40]. Such explanations are extracted using explainable AI (XAI) techniques, which aim to provide insight into how a model arrived at its decision [21, 50]. In this context, explanations may inadvertently provide attackers with insights to enhance their attacks, posing a fresh obstacle to existing security and privacy measures [22, 23, 47].
@@
In this work, we focus on investigating how example-based explanations, specifically counterfactual (CF) explanations, can be exploited to perform MEA within MLaaS scenarios. We precisely focus on CFs due to the unique insights they provide, as they reveal how to minimally alter original data instances to achieve a different model’s outcome, allowing users to interpret the dynamics underlying the model’s prediction shift from the decision boundary [15]. However, due to the proximity of CFs to the decision boundary and their representativeness of the training data, CFs possess a dual nature, serving not only as interpretative tools but also as potential assets for enhancing attacks as they provide additional insights and knowledge for potentially exploiting the vulnerabilities of the model under attack. Specifically, we propose a novel approach based on the concept of knowledge distillation (KD) to perform high-fidelity MEA with CFs as the representation of data, and derive a threat model that have similar functionality of the original model being targeted  [^1] Link to code: https://github.com/FatimaEzzedinee/Knowledge-Distillation-Based-Model-Extraction-Attack-using-GAN-based-Private-CFs. The unique capabilities of KD allow us to handle the problem not just as a standard classification task but as an estimation of the output probability distribution. Moreover, we propose a mitigation approach that employs differential privacy (DP) within the CF generator training pipeline, aiming to reduce the risk associated with providing CF explanations. Our analysis examines the attack performance of our proposed method compared to baselines. Additionally, we assess the impact of DP on the quality of CF and the performance of the MEA. The contribution of this paper is summarized as follows:
@@
  • KD-Based MEA with CFs: This research investigates potential vulnerabilities in MLaaS, with a specific focus on MEA facilitated by KD as extraction techniques and CFs as representatives of the training set. The research simulates an adversarial scenario wherein an attacker employs our proposed KD techniques to extract a substitute for a target model.
@@ 2 Related work
Several works have proposed strategies for performing MEA for classification. Authors in [57] perform successful MEA on different ML models like decision trees, Support Vector Machines, and DNNs by using equation-solving and path-finding algorithms and learning theory. In [45] authors perform MEA on DNNs using active learning with unannotated public data. In [28], authors perform MEA in natural language processing on bert-based APIs, in which they explore transfer learning for MEA. Other works have focused on proposing querying strategies for MLaaS to effectively query the target model and extract accurate insights in addition to performing MEA [44, 26, 42]. In [44] authors train a knockoff network with queried image predictions and propose a reinforcement learning approach that additionally improves query sample efficiency in certain settings and provides performance gains. In [26] authors propose a new method that generates synthetic queries and optimizes training hyperparameters, and then propose a method to detect generic and effective detection of DNN MEA. In [42] authors investigate the type and amount of internal information about the black-box model that can be extracted from querying the deployed model, such as the architecture, optimization procedure, or training data. Authors in [24] improve the query efficiency of attacks by designing learning-based methods and focusing on the real accuracy of the extracted model. Also [29], discuss the tension between the right to explanation and the right to be forgotten for privacy concerns.
@@
Similar to these works, our work aligns with exploring how CF explanations can be exploited for performing MEA. However, our contribution lies in modeling a novel approach to MEAs that exploits more effectively the insights that CFs carry with respect to other approaches that directly train a substitute model on extracted CFs. Moreover, in contrast to previous studies [12, 60], our work presumes that the attacker possesses no prior knowledge of the training set, leading them to query the model with zero knowledge, and propose a novel approach for performing the MEA.
@@
In terms of mitigation strategies, few works have focused on proposing methodologies to generate explanations while mitigating the revelation of sensitive insights about the decision boundary, the training set, or the model architectures. For instance, authors in [46] propose employing DP algorithms to construct feature-based model explanations, while authors in [62] propose an approach to generate differentially private CFs via functional mechanisms. [48] proposes an approach for generating recourse paths leveraging differentially private clustering and demonstrates that constructing a graph on the cluster centers provides private recourse paths as CFs.
@@
Also, our work also exploits DP, however, by integrating it into the CF generation process, and eliminates the necessity for a distinct method to generate private CFs. We leverage the strength of generative adversarial networks (GANs)-based generators in producing high-quality, yet private, CFs. To the best of our knowledge, our work is the first to adapt DP within CF generative-based methods. We specifically focus on GANs-based CF generators due to their ability to generate high-quality CFs compared to other CF generation methods. We show the robustness of the DP-based proposed method against MEA, and we provide a detailed analysis that sheds light on the impact of CFs in exposing the distribution of the training set through well-defined metrics.
3 Scenario and problem formulation
4 Methodology
4.1 Knowledge distillation-based model extraction attack
As an attacker, specifically, we query the model with random queries  [^2] The attacker does not have previous knowledge of the training set. and collect the query output of the target model, and the CFs explaining the output of the target model (Step 3 in Fig. 2). After successfully collecting a set of CFs (Step 4 in Fig. 2), the attacker employs our KD-based MEA approach on the collected dataset as input data to train $t_{\Upsilon}$. The rationale for selecting KD as the extraction method lies in its unique capability to address the problem not only as a standard supervised classification task but also in its proficiency in precisely imitating the probability distribution. Specifically, we propose to train $t_{\Upsilon}$ by minimizing the loss constituted by the threat model classification loss in addition to the distillation loss. To emphasize the importance of mimicking the output probabilistic distribution from the target model to the threat model, we set the distillation loss to the Jensen-Shannon (JS) (1) metric. The choice of JS is to tackle issues related to divergence and ensure a symmetric output, so we suppose that the attacker replaces the KL divergence metric with JS. Afterward, the attacker trains the threat model by minimizing the loss of KD (Step 5 in Fig. 2) until the convergence of the agreement on a separate validation set.
4.2 Private CFs generation with DP
We are now aligned with the service provider and are presenting a strategy to mitigate MEA by avoiding providing Private CFs. Our goal is to examine the effects of incorporating DP into the CF generation process on MEA performance and the quality of the provided CFs considering metrics such as actionability, prediction gain, and realism, which are explained later in sec 5.5. Additionally, we aim to demonstrate how private CFs align with the distribution of the training set. We propose an approach to integrating DP into the CF generation process and generating private CFs. Our methodology does not memorize or expose statistical information about the training set. We propose to incorporate DP into the generator of the CounteRGAN. The primary objective is to prevent the generation of CFs that closely resemble the private training data, thereby reducing the resemblance of CFs with the training data.
5 Experimental settings
5.1 Baseline scenarios
We compare our proposed approach for performing MEA, KD-based MEA, to a baseline approach, presented in [12], and referred to as Direct Train, which trains a threat model as a standard supervised classification task.
We examine three scenarios pertaining to the MLaaS provider’s inclusion or exclusion of CF explanations alongside the predictions of the ML model, 1) No CF: In this scenario, the MLaaS does not provide the user with any CFs. 2) CF: In this case, the MLaaS provides CFs suing counterGAN. 3) Private CF: The MLaaS offers private CFs by implementing our proposed approach for generating private CFs, discussed in Sec. 4.
Applying each MEA approach to the three scenarios results in the following six scenarios: 1) KD-based Private CF. We employ our proposed KD-based approach on CFs extracted from the differentially private generator, and the focus shifts to examining the potential effect of integrating DP in the CF generation. 2) KD-based CF. We employ our proposed KD-based approach, trained with CFs, to understand the influence of KD when applied alongside CF. 3) KD-based No CF. We employ our proposed KD-based approach on the randomly generated data points, and the focus shifts to examining the potential enhancement when KD is utilized independently of CFs and analyzing the power of CFs alongside. 4) Direct No CF. The MEA is performed following the baseline [12]. We train a model directly on the randomly generated data points, and the focus shifts to examining the potential of direct queries and analyzing the power of KD and CFs. 5) Direct CF. We train a model directly on CFs, this dimension is dedicated to evaluating the standalone power of CFs without the incorporation of KD, providing insights into the capabilities of CFs separately. 6) Direct Private CF. We train a model directly on CFs extracted from the differentially private generator, the focus shifts to examining the potential of direct queries and to analyzing the power of KD and CFs alongside.
5.2 Target model
We adhere to consistent training procedures for the target model (to be deployed and queried by the API). We train a DNN comprising of 16 hidden layers to create a complex model. The layer configurations consist of 64, 32, 16, 32, 64, 128, 64, 32, 128, 64, 128, 64, 128, 64, 32, and 16 neurons in each layer, respectively. The activation function is set to gelu [20] for all layers, while the output layer employs softmax to produce model outputs as confidence score probabilities. We specify the optimizer as the Adam optimizer to minimize the cross-entropy loss function, with default parameter initialization provided by Keras. Each target model is trained using 80% of the training dataset  [^3] Experiments were computed with a machine of intel Core i7, a GPU of RTX 3070, and 8 GB of RAM..
5.3 Threat model
5.4 Counterfactual generator
5.5 Evaluation metrics
Effectiveness of MEA: A commonly employed metric for assessing the effectiveness of MEA is the Agreement measure. Agreement measures the degree of alignment between two models, i.e., the similarity in the predictions between the two different ML models. In the context of MEA methods, agreements assess the output similarity between the predictions of an extracted model $t_{\Upsilon}$ and those of the target model $f_{\theta}$ for a given set of data records (see Eqn. 2). Since our goal is to replicate the behavior of a target model with an extracted model, a higher agreement means a more successful MEA, and hence, a more effective strategy for MEAs.
The Realism is a metric used to quantify how closely a data instance fits a known data distribution. Since in our work, we employ DP to avoid providing CF that may expose statistical information about the training set, we employ realism to quantify how well CFs and private CFs fit in with the known data distribution. Furthermore, realism is employed to compare a data point with its associated CF, to assess the authenticity of CFs and to examine the impact of the data perturbation with DP on CF. Inspired by the strategies proposed in [37, 25], we train a denoising autoencoder on the noised training set, and compute the realism of a data point as the reconstruction error as the mean squared error (MSE) of the autoencoder (Eq. 3). A lower value of $realism$ corresponds to a better fit of the data point within the data distribution.
Actionability is a commonly used metric that allows to measure the quality of a CF. Actionability assumes both the number of altered features and the degree of those changes (amount of modification in a feature) in a CF relative to the original data instance. We compute Actionability by taking the L1 norm of the absolute difference between the data point and its corresponding CF (Eq. 4. A low value of actionability (desired) indicates that a relatively small subset of the input features has been perturbed, with respect to the original data instance, to achieve a different outcome by the classifier, and hence, a low actionability suggests better quality CFs.
By computing this comprehensive set of metrics, we aim to explore how DP impacts the feasibility of taking actionable steps and how this influence contributes to the prediction gain outcomes.
6 Results and evaluation
(a) GMSC Dataset
50100200300500100060606060707070708080808090909090Queries NumberKD-based No CFDirect No CF
(b) Credit Card Fraud Dataset
501002003005001000606060607070707080808080Queries NumberKD-based Private CFDirect Private CF
(c) Housing Dataset
Figure 3: The agreement values achieved by the MEA approach in the various scenarios across the a) GMSC dataset, b) Credit Fraud Dataset, c) Housing Dataset with respect to number of queries made.
Figure 3 shows the agreement achieved by KD-based MEA and Direct in the various scenarios with respect to the number of queries made to the API across the GMSC, the Credit Card Fraud, and the California Housing datasets. Results show a consistent pattern across all scenarios as agreement initially increases significantly as the number of queries increases until a point where additional queries cease to yield a notable increase in agreement. This suggests that there is a saturation point in the MEA, and further queries may not necessarily contribute to enhancing the agreement.
We now focus on analyzing the effectiveness of exploiting CFs for performing MEAs, comparing the agreement achieved by each of KD-based MEA and Direct when CF explanations are used to that when only data points are used (i.e., comparing Direct No CF to Direct-CF, and KD-Based No CF to KD-based CF). In both cases, when CFs are exploited, the MEA is more effective, achieving an agreement significantly higher than its counterpart. For instance, in GMSC dataset, Direct-CF achieves an agreement ranging between 80 and 85 while Direct No CF achieves an agreement ranging between 55 and 75 (showing an additional agreement of up to 25%). For Credit Card Fraud dataset, the difference is significantly larger as Direct-CF shows an agreement ranging between 83 and 90 while Direct-no-CF has an agreement ranging between 58 and 75. This difference is also observed for California Housing dataset where exploiting CFs allows for an additional agreement of around 10%. Similarly, analyzing the agreement achieved when employing our proposed approach with CFs to its counterpart (i.e., KD-based CF vs. KD-Based No CF) across the three data sets, we notice that exploiting CFs permits obtaining up to additional 27% of agreement (for GMSC dataset at 50 queries) and, in the worst case, an additional 4% on agreement (for Housing dataset at 1000 queries). These results confirm previous findings in literature (e.g., [12] and [60]) which show the impact of using model’s CF explanations for performing MEA.
In the case when CFs are provided by the MLaaS and are then exploited by the attacker, i.e., KD-based CF vs. Direct CF, results show that across all the 3 datasets, the agreement achieve by KD-based CF agreement always surpasses the Direct-CF, independent of the number of queries. In the case KD-based CF maintains an additional 2% to 6% agreement across all datasets. Similarly, in the case where the MLaaS provides private CFs, i.e., comparing KD-based Private CF to Direct Private CF, KD-based Private CF maintains a slightly additional range from 1% to 2% in most query numbers across all datasets. The observed decrease in agreement of MEA with private CFs may be attributed to the privacy-preserving nature of DP. Differential privacy introduces intentional noise into the data to protect individual privacy, making it more challenging for attackers to accurately extract sensitive information. In the context of KD-based MEA, when DP is employed, the noise introduced during the training process to achieve privacy guarantees could interfere with the teacher-student model’s ability to transfer knowledge much more effectively.
In this consideration, using CFs significantly reduces the number of instances required for MEA to reach a specific high agreement level. For example, on the GMSC dataset, achieving 80% agreement with CFs took just 50 queries, while accomplishing the same level without CFs required over 1000 queries. This pattern holds across datasets, such as the Credit Card Fraud Dataset, where 83% agreement needed only 50 CF queries compared to over 1000 queries without CFs. In essence, the use of CFs consistently leads to a substantial reduction in the number of queries needed to achieve high agreement levels across diverse datasets. In addition to that, our proposed methodology KD-based MEA consistently results in higher agreement levels of MEA compared to the baseline approach. This holds in various scenarios and is independent of whether the MLaaS provides users with CF explanations, and across diverse datasets and query numbers. This analysis emphasizes the effectiveness of our proposed method, utilizing KD to achieve superior agreement levels. Importantly, when CFs are employed, MEA requires significantly fewer instances to reach a specific high agreement level. This highlights the efficiency of incorporating CFs in conjunction with KD to enhance the performance of the MEA.
We now analyze the impact of generating private CFs through incorporating DP in the explainer, i.e., by incorporating DP during the generation process of explanations, on mitigating MEA. Specifically, we aim to quantify the extent of protection offered by providing users with private CFs (as opposed to non-private CFs) such as to prevent attackers from leveraging vulnerabilities associated with non-private CFs. To this end, we compare the performance of each of KD-based CF and Direct CF to its private CF counterparts (KD-based Private CF and Direct Private CF, respectively). For the GMSC dataset, results show that for the relatively low number of queries (50 to 200), the difference is vast between each scenario with private CF and its counterpart. For instance, KD-based Private CF achieves an agreement ranging between 60% and 75% while that KD-based CF obtains an agreement between 83% and 87%.
In the case of the GMSC dataset, we observe that the incorporation of DP shows that our proposed strategy can maintain agreement levels comparable to scenarios without CFs and less than when incorporating traditional CFs, with an agreement range of 60% to 83% with KD-based CF and 60% to 81% with direct training, compared to 57% to 79% with KD-based MEA no CF and a range of 57% to 76% with Direct No CF. For the credit card fraud dataset, a similar trend is observed. With KD-based Private CF, we achieve a range of 63% to 82% and 52% to 79% for a direct-DP train, compared to KD-based no CF of 60% to 81% and a range of 58% to 75% with Direct No CF. For the California housing dataset, the agreement of KD-based CF ranges from 59% to 68% with KD and 57% to 67% with Direct Private CF, compared to 59% to 75% with KD-based no CF, and direct no CF ranges from 57% to 74%. It is also worth noting that, KD-based Private CFis slightly better than direct training over all datasets.
Results suggest that the incorporation of DP can play a crucial role in maintaining agreement levels comparable to scenarios without CFs, as they show that across the three datasets a mitigation can be achieved. In the appendix, we add further discussion about the impact of DP on the generation of CF and discuss what could be the potential reasons behind such results.
7 Conclusion
Appendix A Appendix / supplemental material
A.1 Preliminaries